Calculate the price on its issue date of 100000 face value 90 day commercial paper issued by GE capital Canada if the prevailing market rate of return is 1.932%


1
Expert's answer
2021-05-24T13:20:56-0400

"P=M(1+it)^{-1}"

where;

p=market price

i=market rate

t=time to maturity

"P=100000(1+\frac{1.932}{100}\times \frac{90}{365})^{-1}\\=100000\times0.99525875\\=99525.87509"